OpenStudy (anonymous):

What is the solution to this equation? 3(4x + 3) = 2x - 5(3 - x) + 2 A. x=-22over5 B. x=-4over5 C. x=-4over15 D. x=-22over15

OpenStudy (anonymous):

A is the rite answer.

OpenStudy (anonymous):

how did you get it?

OpenStudy (anonymous):

3(4x + 3) = 2x - 5(3 - x) + 2 12x+9=2x -15 +5x +2 12x + 9 = 7x -13 12x - 7x = -13-9 5x =-22 x= -22/5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

my pleasure.